Delph Company uses a job-order costing system and has two manufacturing departments—Molding and Fabrication. The company provided the following estimates at the beginning of the year:

  

Molding Fabrication Total
Machine-hours 34,000 44,000 78,000
Fixed manufacturing overhead costs $ 760,000 $ 230,000 $ 990,000
Variable manufacturing overhead cost per machine-hour $ 6.10 $ 6.10

  

During the year, the company had no beginning or ending inventories and it started, completed, and sold only two jobs—Job D-70 and Job C-200. It provided the following information related to those two jobs:

  

Job D-70: Molding Fabrication Total
Direct materials cost $ 375,000 $ 327,000 $ 702,000
Direct labor cost $ 220,000 $ 170,000 $ 390,000
Machine-hours 24,000 10,000 34,000

  

Job C-200: Molding Fabrication Total
Direct materials cost $ 270,000 $ 250,000 $ 520,000
Direct labor cost $ 170,000 $ 260,000 $ 430,000
Machine-hours 10,000 34,000 44,000

Delph had no underapplied or overapplied manufacturing overhead during the year.

Required:

1. Assume Delph uses a plantwide predetermined overhead rate based on machine-hours.

a. Compute the plantwide predetermined overhead rate.

b. Compute the total manufacturing cost assigned to Job D-70 and Job C-200.

c. If Delph establishes bid prices that are 120% of total manufacturing cost, what bid prices would it have established for Job D-70 and Job C-200?

d. What is Delph’s cost of goods sold for the year?

1A

Assume Delph uses a plantwide predetermined overhead rate based on machine-hours. Compute the plantwide predetermined overhead rate. (Round your answer to 2 decimal places.)

Predetermined overhead rate per MH

Assume Delph uses a plantwide predetermined overhead rate based on machine-hours. Compute the total manufacturing cost assigned to Job D-70 and Job C-200. (Round your intermediate calculations to 2 decimal places.)

Total Manufacturing Cost
Job D-70
Job C-200

Assume Delph uses a plantwide predetermined overhead rate based on machine-hours. If Delph establishes bid prices that are 120% of total manufacturing cost, what bid prices would it have established for Job D-70 and Job C-200? (Round your intermediate calculations to 2 decimal places.)

Bid price
Job D-70
Job C-200

Assume Delph uses a plantwide predetermined overhead rate based on machine-hours. What is Delph’s cost of goods sold for the year? (Round your intermediate calculations to 2 decimal places.)

Cost of goods sold
Answer #1

1a) Plantwide predetermine overhead rate = (78000*6.1+990000)/78000 = 18.79 per hour

1b) Calculate total manufacturing cost

Job D-70 Job C-200
Direct material 702000 520000
Direct labor 390000 430000
Overhead 34000*18.79 = 638860 44000*18.79 = 826760
Total manufacturing cost 1730860 1776760

1c) Bid price

Job D-70 = 1730860*1.20 = 2077032

Job C-200 = 1776760*1.20 = 2132112

1d) Cost of goods sold = 1730860+1776760 = 3507620